OpenStudy (anonymous):

Solve: (x^2-1)(x^2+3x+2)=0 A)x=-1; x=1; x=2 B)x=-1; x=1; x=-2 C)x=-1; x=0; x=2 D)x=1; x=0; x=2

OpenStudy (anonymous):

factor both polynomials (x^2-1) = (x+1)(x-1) (x^2+3x+2) = (x+2)(x+1) so we get (x+1)(x-1)(x+2)(x+1) = 0

OpenStudy (anonymous):

so that means x = -2, -1, and 1

OpenStudy (anonymous):

@robz8 is correct

OpenStudy (anonymous):

thank you for also showing how you did it, i struggle with these.